* The president of Hill Enterprises. Terri Hill, projects the firm's aggregate demand requirements over the next 6 month as follows:

January 1,600
February 2,500
March 1,800
April 2,000
May 1,900
June 2,200

Her operation manager is considering a new plan. Stockout cost of lost sales is $100.
Develope a constant work force model for this problem that will produce the average demand of this 6 months period in each month. Determine the total cost of this plan.

Her operations manager is considering a new plan. Stock cost of lost sales is $100 per unit. Inventory holding cost is $20 per unit per month. Ignore any idle-time costs.
Answer #1

The aggregate plan is following:

1 Hill Enterprise Ending Production Units Stockout Units 2 Period Month Demand Cell Formula E4 MAX(0,C4-F3-D4 F4 MAX(0,F3-C4+D4 F14 I-E10% E 11 15 F10 F11 F16 -SUM(F14:F15 Inventor Copy to E4:E9 F4:F9 0 December 1 January 1,600 2,000 2,000 2,000 2,000 2,000 2,000 0 100 0 0 0 0 100 100 52Fe2,500 6 3 March 0 200 200 rua 1,800 2,000 1,900 2,200 4 April 8 5 May 9 6 June 10 Total 11 Cost/Category 12 100 1,200 S20 The total stockout cost of lost sales 10,000 The total inventory carrying cost 24,000 The total cost, excluding normal time labor costs34,000 15 17

EXCEL FORMULAS:

Cell Formula Copy to
E4 =MAX(0,C4-F3-D4) E4:E9
F4 =MAX(0,F3-C4+D4) F4:F9
F14 =E10*E11
F15 =F10*F11
F16 =SUM(F14:F15)

Total cost of this plan = $ 34,000